achung6104:

For ΔABC, ∠A = 8x - 10, ∠B = 10x - 40, and ∠C = 3x 20. If ΔABC undergoes a dilation by a scale factor of 1 2 to create ΔA'B'C' with ∠A' = 6x 10, ∠B' = 70 - x, and ∠C' = 10x 2 , which confirms that ΔABC∼ΔA'B'C by the AA criterion?

russkaya05:

I think its C. C) ∠B = ∠B' = 33° and ∠C = ∠C' = 80° basically the triangle has a total of 180 degrees. In order to get the x: 180 = 4x + 7 + 2x + 3 + 6x -10 180 = 12x - 0 x = 15 ∠A = 4x + 7 = 67 ∠B = 2x + 3 = 33 ∠C = 6x - 10 = 80 hopefully this is what you were looking for